Suppose a firm’s business operations mirror movements in the economy as a whole very closely—that is, the firm’s asset beta is 1. Find the equity beta for this firm for debt-equity ratios of 0, 1.8, 6.4, and 27. (Do not round intermediate calculations and round your answers to 1 decimal place, e.g., 32.1.)

Debt-Equity Ratio Equity Beta
0
1.8
6.4
27
Answer #1
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 1*(1+((1-0)*(0)))
levered beta = 1
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 1*(1+((1-0)*(1.8)))
levered beta = 2.8
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 1*(1+((1-0)*(6.4)))
levered beta = 7.4
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 1*(1+((1-0)*(27)))
levered beta = 28